OpenStudy (anonymous):

Which expression represents the area of the triangle? base=2x+12, height x+7 A) 3x + 19 B) 3x2 + 84 C) x2 + 13x + 42 D) 2x2 + 26x + 84

OpenStudy (anonymous):

A = bh/2 So, you would multiply the 2 expression for "b" and "h" (that are in "x") and divide by "2". Hint: It's easier to divide the "base" (or b) expression first by "2" and then simply multiply that by the expression for "h".

OpenStudy (anonymous):

So, what is the expression for for the base divided by "2"?

OpenStudy (anonymous):

1x+6

OpenStudy (anonymous):

Very good. So, now, all you have to do is multiply your answer by "x + 7" (x + 6)(x + 7) = ?

OpenStudy (anonymous):

Are you stuck?

OpenStudy (anonymous):

46

OpenStudy (anonymous):

When you have an expression like: (x + a)(x + b) It is: x(x + b) + a(x + b) = x^2 + bx + ax + ab = x^2 + (a + b)x + ab So, just substitute a = 6 and b = 7 in that last equation

OpenStudy (anonymous):

thank you this is so confusing

OpenStudy (anonymous):

Math can be, but if you take it step by step, it won't trick you. IMath is like a friend that speaks a foreign language.

OpenStudy (anonymous):

i'll practice and see if i can learn the langage

OpenStudy (anonymous):

That's a great attitude! You'll succeed if you stick with it. And you might want to go over that equality expansion a couple of times because it could clear things up for you.

OpenStudy (anonymous):

ok thanks so much

OpenStudy (anonymous):

You're very welcome and nice working with you!

OpenStudy (anonymous):

And thx for the recognition! @TestTakingLifesaver Good luck in all of your studies.